OpenStudy (anonymous):

Ferris purchased a car for $16,720. He made a down payment of $1,940. He applied for a five-year installment loan with an interest rate of 8.1%. What is the total cost of the car after five years?

OpenStudy (anonymous):

compount interest formula: A=P(1+r100)n A = amount after "n" years = to find P = amount for interest = 16720-1940 r = APR = 8.1 n = number of years = 5